OpenStudy (anonymous):

Solve for H: A = L • W • H please help have no idea how to do this lol ^...^

OpenStudy (akashdeepdeb):

A = LWH Now we have to get the value of 'H' in this. So we have to get rid of the 'LW' on the Right Hand side. Now as 'LW' is multiplied to 'H' to get RID of it, we need to divide it by 'LW' So A = LWH \[\frac{A}{LW} = \frac{LWH}{LW}\] Thus \[H = \frac{A}{LW}\] Understood? :)

OpenStudy (anonymous):

yes thankxz a bunch

OpenStudy (akashdeepdeb):

:)